-- RobertaCarini - 27 Feb 2019

INAF queue observing, 2019 February 27 - 28

Observer: R. Carini, A. Rossi
Telescope Operator: J. Williams
Support Astronomer: D. Thompson
AO Support D. Miller
Instrument: LUCI1,2, AO

Summary

At the sun set some clouds on the sky

Night duration (12 degree el twilight) 2:00- 13:00 UT = 11 h 00m

Observing time: 10h 40min (97%)

Weather time loss: h m (%)

technical time loss 0h 20m (3%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Severgnini PSF A5 UCAC 589-042207 LUCI2-AO K 3 m luci2.0029;luci2.0031-0032  
Severgnini A5 SDSS J075223 LUCI2-AO K 0.3h luci2.0033-0053  
Severgnini PSF A5 UCAC 589-042207 LUCI2-AO K 3m luci2.0054-0056 completed
Severgnini PSF A1 NOMAD 1543-0162089 LUCI2-AO K 3m luci2.0057-0059  
Severgnini A1 SDSS J085512 LUCI2-AO K 0.2h luci2.0060-0064 completed
Severgnini PSF A1 NOMAD 1543-0162089 LUCI2-AO K 3m luci2.0065-0067  
Gilli telluric HIP56736 LUCI1,2 HKspec, ls   luci1.0015-0018, luci2.0076-0079  
Gilli L SDSS J1030+0524 LUCI1,2 HKspec, ls 2h+2h luci1.0024-0059, luci2.0085-0120 completed
Severgnini PSF A4 NOMAD 1010-0202834 LUCI2-AO K 3m luci2.0124-0126  
Severgnini A4 SDSS J115548 LUCI2AO K 0.6 luci2.0127-0156 completed
Severgnini PSF A4 NOMAD 1010-0202834 LUCI2-AO K 3m luci2.0157-0159  
Severgnini PSF A6 NOMAD 1391-0251115 LUCI2-AO K 3m luci2.0161-0163  
Severgnini A6 SDSS J145110 LUCI2-AO K   lucl2.0164-0193 completed
Severgnini PSF A6 NOMAD 1391-0251115 LUCI2-AO K 3m luci2.0194-0196  
Nardini q9 SDSS J142543 LUCI1 G150, Ks 0.67h luci1.0064-0084 completed
telluric q9   LUCI1 Ks luci1.0088-0089    
calibration flat + arc Gilli, Severgnini, Nardini LUCI1, LUCI2        

Details

01:30 AO check

02:15 the sky is clear, we can start with AO program of Severgnini, target A5 (SDSS J075223.35+273643.1)

02:24 send the psf star for A5, UCAC 589-042207
  • the Field Stop for N30 Camera is not aligned. we do it but closing the loop takes a lot of time
  • 02:40 Field Stop Alignment for N30 camera luci2.20190228.0001-0027

02:45 send again the script for the psf star UCAC 589-042207 luci2.0028-0029
  • seeing on dimm 1.00"
  • in the image luci2.0028 the psf star is double, we do not know why
  • seeing on dimm 0.9", fwhm 0.7"

02:49 send the script of the psf star of A5 == luci2.0030-0032 luci2.0030 is not good *seeing on dimm 0.9"

03:07 send the script of the sceintific starget A5 luci2.0033-0053
  • seeing on dimm 0.9"; fwhm 0.9", estimated magnitude 16.4
  • 03:27 seeing 0.95"on dimm, fwhm 0.95".
  • fwhm on image 9-10pixels~ 0.15" measured on stars appearing on the bottom border every 2 offsets
  • 03:35 loop open! Doug closes the loop manually 1-2 min lost.
  • 03:45 seeing on dimm 1.0", mag 16.5, binning 4, fwhm on image 0.15"

03:48 send the script of the psf star of A5 luci2.0054-0056
  • seeing on dimm and fwhm 0.95".

03:58 slew to PSF for the target A1 (J085512), NOMAD 1543-0162089, luci2.0057-0059
  • seeing on dimm 1.0"-1.1"

04:10 target A1
  • 04:12 re-set the correction for the astigmatism.
  • 04:20 start the science integration luci2.0060-0064

04:20 PSF for the target A1 luci2.0065-0067
  • seeing on dimm 0.9"-1.1"

04:45 PSF for the target A3 J091110, NOMAD 1041-0170214;
  • no star find in the guider but the scripts seems correct
  • 04:57 send again the script but it doesn't work, the TO does the collimation again with a brighter star.
  • 05:05 send again the script but there is something wrong with the guide star, maybe it is a galaxy.

05:08 slew to Gilli program, telluric star for target L
  • collimation preset
  • the script started by itself, error in the mask with LUCI1 and wiht the filter in LUCI2, the SA fixes the issues.
  • 05:34 send again the script
  • acq luci1.0012-0014, luci2.0073-0075
  • 05:42 science luci1.0015-0018, luci2.0076-0079

05:46 slew to Gilli target L, new OB
  • seeing 1" on DIMM
  • acq luci1.0019-0023, luci2.0080-0084
  • we adjusted manually the offsets.
  • the script had a pause after the blind offset and was not going on.
  • 06:05 science luci1.0024-0041, luci2.0085-0102
  • seeing 1.2" on DIMM
  • 06:28 fwhm 0.9",
  • 06:32 seeing on dimm 0.83"
  • 07:30 send again the script luci1.0042-0059, luci2.0103=0120
  • seeing on dimm 0.8"
  • we have realized that the script is using readout LIR instead of recommended MER mode.
  • seeing 0.7" at 08:05 on DIMM
  • 08:40 seeing on dimm 0.8"

08:45 we slew to Severgnini Target A4, PSF star the field stop is mis-slignemnt, we do the alignment luci2.0124-0126
  • 09:12 take the images of PSF star luci2.0124-0126

09:20 send the science script TARGET A4
  • 09:28 loop closed, science starts. luci2.0127-0156
  • seeing on dimm 0.77", fwhm 0.63"

10:12 PSF star, Target A4 luci2.0157-0159

10:25 slew to PSF for Target A6 luci2.0161-0163 * luci2.0160 check the alignment of the field stop, it was ok

10:43 Target A6 lucl2.0164-0193
  • 10:47 seeing on dimm 1.05", fwhm 0.74"
  • 10:55 seeing on dimm 0.9"-1.0", fwhm 0.7"

11:25 slew to PSF for Target A6
  • comunication issues
  • 11:42 send the preset again luci2.0194-0196

11:55 slew to Nardini's program, target q9
  • TO authorizes LUCI1 only
  • 12:02 send the script acq luci1.0060-0063
  • sci luci1.0064-0084

1:06 telluric for q9, HIP 65280
  • acq luci1.0085-0087
  • 01:14 sci luci1.0088-0089
  • lost guide star

CALIBRATIONS Gilli
  • luci1.0090-0094, luci2.0197-0201 flush
  • luci1.0095-0099, luci2-0202-0206 flat, halo2
  • luci1.0100-0101, luci2.0207-0208 arc Ne
  • luci1.0102-0103, luci2.0209-0210 arc Ar
  • luci1.0104-0105, luci2.0211-0212 Xe

Severgnini
  • luci2.0213-0217 flush
  • luci2.0218-0222 flat dome K

Nardini
  • luci1.0106-0110 flush
  • luci1.0111-0115 flat
  • luci1.0116-0117 arc Ne
  • luci1.0018-0019 arc Ar
  • luci1.0020-0021 arc Xe
Topic revision: r13 - 04 Mar 2019, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback